OpenStudy (whtcutie):

Solve the expression p^2+4p=1 by completing the square. show work. how do I do that? need help please!

OpenStudy (lgnd_88):

p = 1/2 = 0.500

OpenStudy (whtcutie):

how did you get that? @LGND_88

OpenStudy (mathmale):

"Solve the expression p^2+4p=1 by completing the square." Actually, this is an equation, not an expression. You are told to use a particular method (completing the square) to answer this. Thus, shelve the quadratic formula and factoring for now. Solve the expression p^2+4p=1 by completing the square. Start with the first two terms, both of which involve the variable, p.

OpenStudy (mathmale):

4 is the coefficient of the 'p' term. Take half of this coefficient. That gives you 2. Square this result. What do you get? Now do this: Start with p^2 + 4p = 1. Following the "4p" term, write the square you found above, adding it to p^2+4p, and then subtracting it from p^2 +4p+4. What do you get? As soon as I have your answers, we'll continue.

OpenStudy (whtcutie):

wait, so i do p^2+2p=1 hats for the first thing you want me to do right?

OpenStudy (whtcutie):

@3mar can you maybe help me with the new one that I posted?

OpenStudy (3mar):

Well, I am here. I will not be late for any help.

OpenStudy (mathmale):

Suggest you work on one question at a time. "Start with p^2 + 4p = 1. Following the "4p" term, write the square you found above, adding it to p^2+4p, and then subtracting it from p^2 +4p+4. What do you get?" Yes, actually write out p^2 +4p =1 Follow my instructions, please. If you're unsure of what you want to do, please ask questions.

OpenStudy (whtcutie):

solve the equation p^2+4p=1 by completing the square. show work. @3mar

OpenStudy (whtcutie):

@mathmale ok yes but I am getting confuse.

OpenStudy (mathmale):

Is it really helpful to you, or to me, if you say "I'm getting confuse?" I asked you to ask questions to clear up whatever it is that confuses you. Please go back and read my suggestions. If they're not clear, tell me where they're not clear and what you want to know.

OpenStudy (3mar):

I will simplify it for you In Sha' Allah! What do you need to make \(p^2+4p\) equal to \(p^2+4p+4\)???

OpenStudy (mathmale):

Actually, we need to "make" p^2+4p equal to p^2+4p + 4 - 4. This follows my suggestions, above.

OpenStudy (mathmale):

You begin with p^2 + 4p = 1 and rewrite this as p^2 + 4p = 1 (as I indicated earlier) I ask you to take the coefficient of the p term (it is 4), halve that and square your result. Can you do that?

OpenStudy (whtcutie):

would you need to make 4p into a square root? so it would be 2p?

OpenStudy (mathmale):

cutie, I ask you to take the coefficient (4) of the 4p term, take half of that coefficient (which would be 4/2), and then to square this result. What is the square of (4/2)? Same question: What is the square of 2?

OpenStudy (mathmale):

No, do not try to "make 4p into a square root."

OpenStudy (3mar):

I think this would help you. I am with you if you faced any difficulties!

OpenStudy (mathmale):

There simply isn't any way to make this "complete the square" process easier. Once you've used it a few times, it'll seem easy. What is the square of 2? 2^2=?

OpenStudy (whtcutie):

@3mar ok thank you that helped so the square of 2 is 4

OpenStudy (3mar):

No problem. I am here for your help. Yes that is right \(2^2=4\)!

OpenStudy (whtcutie):